# Dirac Fermions and Topological Phases in Magnetic Topological Insulator Films

### v2

@@ A.1 Symmetry analysis of solutions
There exists in fact a hidden symmetry in the model, namely, the mirror symmetry about the $x$-$y$ plane. Effectively, it will also bring $z$ to $-z$ as an inversion, but with an additional operation that rotates spin angular momentum by $\pi$ phase, i.e., such a $z$-Mirror symmetry $\mathcal{M}_{z}$ is a combination of $\mathcal{P}_{z}$ and a $\mathcal{C}_{2z}$ rotation, which then acts on the basis as
@@
$\Phi_{++}$ $=\Phi_{1},\leavevmode\nobreak\ \Phi_{+-}=\Phi_{2},$ (176)
$\Phi_{-+}$ $=\Phi_{3},\leavevmode\nobreak\ \Phi_{--}=\Phi_{4}.$
@@ A.2 Equivalent block Hamiltonian
The projection procedure works under the given basis representation $H_{TI}(\bm{k})$, which is formally $H=\bra{\Phi}H_{TI}(\bm{k})\ket{\Phi}$, with
@@
where the integral is done from $-L/2$ to $L/2$. Clearly, projection on $H_{1d}$ would give $\mathrm{diag}(m,-m,-m,m)$, then we only need to deal with $H_{\parallel}(\bm{k})=\lambda_{\parallel}(\bm{k}\cdot\bm{\sigma})\tau_{x}=\lambda_{\parallel}(k_{x}\sigma_{x}+k_{y}\sigma_{y})\tau_{x}$ term. Since $H_{\parallel}(\bm{k})$ is purely off-diagonal, it is easy to conclude that
$\braket{\Phi^{n}_{i}}{H_{\parallel}}{\Phi^{n^{\prime}}_{i}}$ $=0,\ i=1,2,3,4$
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{3}}$ $=0=\braket{\Phi^{n}_{2}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}}.$
@@
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}}$ $=\lambda_{\parallel}k_{-}|C^{n}C^{n^{\prime}}|\int\mathrm{d}z\leavevmode\nobreak\ i\lambda_{\perp}t_{\perp}[\eta^{n}(f^{n}_{+})^{*}f^{n^{\prime}}_{-}+\eta^{n^{\prime}}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}]=0,$
$\braket{\Phi^{n}_{2}}{H_{\parallel}}{\Phi^{n^{\prime}}_{3}}$ $=\lambda_{\parallel}k_{-}|C^{n}C^{n^{\prime}}|\int\mathrm{d}z\leavevmode\nobreak\ i\lambda_{\perp}t_{\perp}[\eta^{n}(f^{n}_{-})^{*}f^{n^{\prime}}_{+}+\eta^{n^{\prime}}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}]=0,$
@@
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{2}}$ $=\int\mathrm{d}z\ \lambda_{\parallel}k_{-}\varphi^{\dagger}(\lambda_{\perp})\tau_{x}\chi(-\lambda_{\perp})=\lambda_{\parallel}k_{-}\delta_{nn^{\prime}},$
$\braket{\Phi^{n}_{3}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}}$ $=\int\mathrm{d}z\ \lambda_{\parallel}k_{-}\varphi^{\dagger}(\lambda_{\perp})\tau_{x}\chi(-\lambda_{\perp})=\lambda_{\parallel}k_{-}\delta_{nn^{\prime}},$
@@
$H(\bm{k})=\int\mathrm{d}z\leavevmode\nobreak\ (U^{c})^{\dagger}(\bm{k},z)H_{TI}(\bm{k},-i\partial_{z})U^{c}(\bm{k},z).$ (181)
@@ A.3 Analytic expression for mass term
where $g^{+}(\xi)=\tan(\xi L/2)/\xi$. We shall assume $\lambda_{\perp}>0,\leavevmode\nobreak\ t_{\perp}>0$ in the following discussion, without losing generality, and $m_{0}(\bm{k})$ controls the expression form.
@@
fulfills the requirement. Notice that $\gamma<1/2$ is assumed, which should not bother the self-consistent solution. Meanwhile, since $\xi_{1}=0$ leads to degenerate eigenvalue ±$\pm\xi_{1}$, then one should generally assume another solution as
@@ A.4 Finite-size correction to mass term
$\bullet$ $40<\gamma<1/$ 4, we choose
@@
considering that $40<\gamma<1/$ 4 in this zone. Again, one obtains
@@
However, as a comment, in numerical calculation, $E$ in zone $40<\gamma<1/$ 4 is suppressed into zero in a much slower manner, which is caused by exponential cancellation between $\sinh$ and $\exp$. Nevertheless, since $\sqrt{1-4\gamma}<1$ in the region, we conclude that the exponential increasing is always slower than the decaying, which finally pushes the state to zero energy for $L\rightarrow+\infty$.
@@
To obtain an effective model, we start from solving $\mathcal{H}_{1d}$ and notice that $[\mathcal{H}_{1d}(\bm{k}),\sigma_{z}]=0$, from which we could let
$\mathcal{H}_{1d}(\bm{k})\zeta_{s}\otimes\ket{\phi^{s}(\bm{k})}=\zeta_{s}\otimes\mathcal{H}^{s}_{1d}(\bm{k})\ket{\phi^{s}(\bm{k})},$ (216)
@@
$\sigma_{z}\zeta_{s}=s\zeta_{s},\leavevmode\nobreak\ s=\pm.$ (217)
@@
$[(t_{\perp}\tau_{z}+i\frac{\lambda_{\perp}}{2}s\tau_{x})\beta^{-1}+(M\tau_{z}-E)+(t_{\perp}\tau_{z}-i\frac{\lambda_{\perp}}{2}s\tau_{x})\beta]\phi=0,$ (220)
@@
$\sin\xi_{\alpha}=\sqrt{1-\cos\xi_{\alpha}^{2}},\leavevmode\nobreak\ \alpha=1,2.$ (239)
Basically, there are three variables $\xi_{1},\xi_{2}$ and $m$, together with three equations above, then it is in a sense some exact system of equations but a non-linear transcendental version. From this set of equations, one may expect $L_{z}$ solutions $m_{n}(\bm{k}),n=1,2,\cdots,L_{z}$ including one surface state and $L_{z}-1$ purely trivial bulk states, if within suitable choice of parameters. And the other set of $L_{z}$ solutions are just chiral partners with $-m_{n}(\bm{k})$. Notice that these $2L_{z}$ solutions compose eigenvalues for one $H_{1d}^{s}$, then by counting $s=\pm$ there are in fact $4L_{z}$ solutions in total, which is expected from the matrix form of $H_{1d}$.
@@
$\Phi^{n}_{1}=\zeta_{+}\otimes\varphi(+)$ $=\begin{pmatrix}\varphi^{n}(+)\\ 0\end{pmatrix},\leavevmode\nobreak\ \Phi^{n}_{2}=\begin{pmatrix}0\\ \chi^{n}(-)\end{pmatrix},$ (241)
$\Phi^{n}_{3}$ $=\begin{pmatrix}\chi^{n}(+)\\ 0\end{pmatrix},\leavevmode\nobreak\ \Phi^{n}_{4}=\begin{pmatrix}0\\ \varphi^{n}(-)\end{pmatrix},$
@@
The basis here shares the same symmetry analysis as within the continuum model, while here the parity and mirror symmetries can be written down explicitly in the off-diagonal matrix form, with $\sigma_{0}\tau_{z}$ and $-i\sigma_{z}\tau_{z}$ as the anti-diagonal elements, respectively. And especially, by combining the mirror and spin-$z$ index, we assign $\Phi_{i}^{n}=\Phi^{n}_{\chi,s}$ with
$\Phi^{n}_{++}$ $=\Phi^{n}_{1},\leavevmode\nobreak\ \Phi^{n}_{+-}=\Phi^{n}_{2},$ (242)
$\Phi^{n}_{-+}$ $=\Phi^{n}_{3},\leavevmode\nobreak\ \Phi^{n}_{--}=\Phi^{n}_{4}.$
@@
[…]
@@
$\braket{\Phi^{n}_{i}}{H_{\parallel}}{\Phi^{n^{\prime}}_{i}}$ $=0,\ i=1,2,3,4,$
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{3}}$ $=0=\braket{\Phi^{n}_{2}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}}.$
@@
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\sum_{l_{z}}|C|^{2}i\lambda_{\perp}t_{\perp}[\eta^{n}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}+\eta^{n^{\prime}}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}]=0,$
$\braket{\Phi^{n}_{2}}{H_{\parallel}}{\Phi^{n^{\prime}}_{3}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\sum_{l_{z}}|C|^{2}i\lambda_{\perp}t_{\perp}[\eta^{n}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}+\eta^{n^{\prime}}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}]=0,$
@@
$\braket{\Phi^{n}_{1}}{H_{\parallel}}{\Phi^{n^{\prime}}_{2}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\delta_{nn^{\prime}}=\braket{\Phi^{n}_{3}}{H_{\parallel}}{\Phi^{n^{\prime}}_{4}},$

### v3

@@ A.1 Symmetry analysis of solutions
There exists in fact a hidden symmetry in the model, namely, the mirror symmetry about the $x$-$y$ plane. Effectively, it will also bring $z$ to z-z as an inversion, but with an additional operation that rotates spin angular momentum by $\pi$ phase, i.e., such a $z$-Mirror symmetry $\mathcal{M}_{z}$ is a combination of $\mathcal{P}_{z}$ and a $\mathcal{C}_{2z}$ rotation, which then acts on the basis as
@@
$\Phi_{++}$ $=\Phi_{1},\penalty 10000\ \Phi_{+-}=\Phi_{2},$ (176)
$\Phi_{-+}$ $=\Phi_{3},\penalty 10000\ \Phi_{--}=\Phi_{4}.$
@@ A.2 Equivalent block Hamiltonian
[…]
@@
where the integral is done from L/2-L/2 to $L/2$. Clearly, projection on $H_{1d}$ would give $\mathrm{diag}(m,-m,-m,m)$, then we only need to deal with $H_{\parallel}(\bm{k})=\lambda_{\parallel}(\bm{k}\cdot\bm{\sigma})\tau_{x}=\lambda_{\parallel}(k_{x}\sigma_{x}+k_{y}\sigma_{y})\tau_{x}$ term. Since $H_{\parallel}(\bm{k})$ is purely off-diagonal, it is easy to conclude that
$\braket{\Phi^{n}_{i}|H_{\parallel}|\Phi^{n^{\prime}}_{i}}$ $=0,\ i=1,2,3,4$
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{3}}$ $=0=\braket{\Phi^{n}_{2}|H_{\parallel}|\Phi^{n^{\prime}}_{4}}.$
@@
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{4}}$ $=\lambda_{\parallel}k_{-}|C^{n}C^{n^{\prime}}|\int\mathrm{d}z\penalty 10000\ i\lambda_{\perp}t_{\perp}[\eta^{n}(f^{n}_{+})^{*}f^{n^{\prime}}_{-}+\eta^{n^{\prime}}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}]=0,$
$\braket{\Phi^{n}_{2}|H_{\parallel}|\Phi^{n^{\prime}}_{3}}$ $=\lambda_{\parallel}k_{-}|C^{n}C^{n^{\prime}}|\int\mathrm{d}z\penalty 10000\ i\lambda_{\perp}t_{\perp}[\eta^{n}(f^{n}_{-})^{*}f^{n^{\prime}}_{+}+\eta^{n^{\prime}}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}]=0,$
@@
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{2}}$ $=\int\mathrm{d}z\ \lambda_{\parallel}k_{-}\varphi^{\dagger}(\lambda_{\perp})\tau_{x}\chi(-\lambda_{\perp})=\lambda_{\parallel}k_{-}\delta_{nn^{\prime}},$
$\braket{\Phi^{n}_{3}|H_{\parallel}|\Phi^{n^{\prime}}_{4}}$ $=\int\mathrm{d}z\ \lambda_{\parallel}k_{-}\varphi^{\dagger}(\lambda_{\perp})\tau_{x}\chi(-\lambda_{\perp})=\lambda_{\parallel}k_{-}\delta_{nn^{\prime}},$
@@
$H(\bm{k})=\int\mathrm{d}z\penalty 10000\ (U^{c})^{\dagger}(\bm{k},z)H_{TI}(\bm{k},-i\partial_{z})U^{c}(\bm{k},z).$ (181)
@@ A.3 Analytic expression for mass term
where $g^{+}(\xi)=\tan(\xi L/2)/\xi$. We shall assume $\lambda_{\perp}>0,\penalty 10000\ t_{\perp}>0$ in the following discussion, without losing generality, and $m_{0}(\bm{k})$ controls the expression form.
@@
fulfills the requirement. Notice that $\gamma<1/2$ is assumed, which should not bother the self-consistent solution. Meanwhile, since $\xi_{1}=0$ leads to degenerate eigenvalue $\pm\xi_{1}$, then one should generally assume another solution as
@@ A.4 Finite-size correction to mass term
$\bullet$ $0<\gamma<1/4$, we choose
@@
considering that $0<\gamma<1/4$ in this zone. Again, one obtains
@@
However, as a comment, in numerical calculation, $E$ in zone $0<\gamma<1/4$ is suppressed into zero in a much slower manner, which is caused by exponential cancellation between $\sinh$ and $\exp$. Nevertheless, since $\sqrt{1-4\gamma}<1$ in the region, we conclude that the exponential increasing is always slower than the decaying, which finally pushes the state to zero energy for $L\rightarrow+\infty$.
@@
To obtain an effective model, we start from solving $\mathcal{H}_{1d}$ and notice that $\mathcal{H}_{1d}(\bm{k}),\sigma_{z}]=0$, from which we could let
[…]
@@
$\sigma_{z}\zeta_{s}=s\zeta_{s},\penalty 10000\ s=\pm.$ (217)
@@
$(t_{\perp}\tau_{z}+i\frac{\lambda_{\perp}}{2}s\tau_{x})\beta^{-1}+(M\tau_{z}-E)+(t_{\perp}\tau_{z}-i\frac{\lambda_{\perp}}{2}s\tau_{x})\beta]\phi=0,$ (220)
@@
$\sin\xi_{\alpha}=\sqrt{1-\cos\xi_{\alpha}^{2}},\penalty 10000\ \alpha=1,2.$ (239)
Basically, there are three variables $\xi_{1},\xi_{2}$ and $m$, together with three equations above, then it is in a sense some exact system of equations but a non-linear transcendental version. From this set of equations, one may expect $L_{z}$ solutions $m_{n}(\bm{k}),n=1,2,\cdots,L_{z}$ including one surface state and $L_{z}-1$ purely trivial bulk states, if within suitable choice of parameters. And the other set of $L_{z}$ solutions are just chiral partners with $m_{n}(\bm{k})$. Notice that these $2L_{z}$ solutions compose eigenvalues for one $H_{1d}^{s}$, then by counting $s=\pm$ there are in fact $4L_{z}$ solutions in total, which is expected from the matrix form of $H_{1d}$.
@@
$\Phi^{n}_{1}=\zeta_{+}\otimes\varphi(+)$ $=\begin{pmatrix}\varphi^{n}(+)\\ 0\end{pmatrix},\penalty 10000\ \Phi^{n}_{2}=\begin{pmatrix}0\\ \chi^{n}(-)\end{pmatrix}$, (241)
$\Phi^{n}_{3}$ $=\begin{pmatrix}\chi^{n}(+)\\ 0\end{pmatrix},\penalty 10000\ \Phi^{n}_{4}=\begin{pmatrix}0\\ \varphi^{n}(-)\end{pmatrix}$,
@@
The basis here shares the same symmetry analysis as within the continuum model, while here the parity and mirror symmetries can be written down explicitly in the off-diagonal matrix form, with $\sigma_{0}\tau_{z}$ and $i\sigma_{z}\tau_{z}$ as the anti-diagonal elements, respectively. And especially, by combining the mirror and spin-$z$ index, we assign $\Phi_{i}^{n}=\Phi^{n}_{\chi,s}$ with
$\Phi^{n}_{++}$ $=\Phi^{n}_{1},\penalty 10000\ \Phi^{n}_{+-}=\Phi^{n}_{2},$ (242)
$\Phi^{n}_{-+}$ $=\Phi^{n}_{3},\penalty 10000\ \Phi^{n}_{--}=\Phi^{n}_{4}.$
@@
$\braket{\Phi|H_{\text{Film}}|\Phi}=\braket{\Phi|H_{1d}|\Phi}+\braket{\Phi|H_{\parallel}|\Phi},$ (243)
@@
$\braket{\Phi^{n}_{i}|H_{\parallel}|\Phi^{n^{\prime}}_{i}}$ $=0,\ i=1,2,3,4,$
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{3}}$ $=0=\braket{\Phi^{n}_{2}|H_{\parallel}|\Phi^{n^{\prime}}_{4}}.$
@@
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{4}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\sum_{l_{z}}|C|^{2}i\lambda_{\perp}t_{\perp}[\eta^{n}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}+\eta^{n^{\prime}}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}]=0,$
$\braket{\Phi^{n}_{2}|H_{\parallel}|\Phi^{n^{\prime}}_{3}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\sum_{l_{z}}|C|^{2}i\lambda_{\perp}t_{\perp}[\eta^{n}(f_{-}^{n})^{*}f_{+}^{n^{\prime}}+\eta^{n^{\prime}}(f_{+}^{n})^{*}f_{-}^{n^{\prime}}]=0,$
@@
$\braket{\Phi^{n}_{1}|H_{\parallel}|\Phi^{n^{\prime}}_{2}}$ $=\lambda_{\parallel}(\sin(k_{x}a)-i\sin(k_{y}b))\delta_{nn^{\prime}}=\braket{\Phi^{n}_{3}|H_{\parallel}|\Phi^{n^{\prime}}_{4}},$